Customer Day Browser Time (min) 12.0 19.5 Pages Viewed Amount Spent (S) Mon Chrome Mon Chrome Wed Chrome Sun Chrome 54.52 94.90 26.68 44.73 66.27 67.80 36.04 55.96 70.94 68.73 54.04 48.05 64.16 158.51 84.12 59.20 91.62 126.40 Wed Other Tue Firefox Sat Firefox 10.5 Fri Firefox Wed Firefox 12.7 24.7 13.3 Tue Chrome Sat Other Sun Firefox Sun Other Mon Firefox 12 13 24.4 Fri Chrome Thu Chrome Thu Chrome Mon Chrome Fri Other Wed Chrome 18 14.0 15.1 10.2 68.45 32.69 78.58 74.43 32.73 20 21 Sat Firefox Tue Chrome Sun Chrome 48.66 54.66 Tue Firefox Fri Firefox Wed Chrome Sat Chrome Sat Firefox Fri Firefox 34.69 27.91 155.30 120.25 16.9 Mon Other Thu Chrome Fri Firefox 33 Sun Chrome Fri Chrome Tue Chrome Mon Firefox 41.20 134.40 37.17 52.09 31 59.99 37 Thu Firefox 55.58 Sat Chrome Tue Chrome Wed Chrome 36.48 67.44 60.14 70.38 110.65 Tue Other 15.9 Sat Chrome Fri Firefox 44 Mon Chrome
45 46 47 48 49 50 Mon Chrome Fri Firefox Thu Chrome Fri Chrome Mon Other Fri Chrome 30.1 13.7 8.1 9.7 7.3 13.4 104.23 68.17 17.84 103.15 52.15 98.75 4
#1.A For the full HeavenlyChocolates data set, fill in the following table (to two decimals): Time Pages Amt Spent Mean Standard Deviation Count Confidence Level(99.0%) #1.B maintain/reject the null hypothesis that the mean value of Time is 12 min, at the .01 level because #1.C maintain/reject the null hypothesis that the mean value of Amt Spent is $70 at the .01 level because #2.A For the full HeavenlyChocolates data set, fill in the correlation table (to two decimals): TimePagesAmt Spent Time Pages Amt Spent #28 which variables, if any, are correlated under the 0.7 rule?- values of Time and Amt Spent by day of the week (to two decimals): Time Amt Spent Sun Mon Tue Wed Thu Fri Sat #3.B The greatest mean value of Time, which is a demand on the firms servers, occurs on which day of the week?
Answer #1


> x=c(12,19.5,8.5,11.4,11.3,10.5,11.4,4.3,12.7,24.7,13.3,14.3,11.7,24.4,8.4,9.6,23.3,14,5.6,15.1,16.3,10.2,8,8,9.6,11,16.9,6,32.9,11.87.1,18,11.8,9.1,7.8,16.5,6.2,11.3,10.6,5.0,15.9,18.1,10.8,13.3,30.1,13.7,8.1,9.7,7.3,13.4)
Error: unexpected numeric constant in "x=c(12,19.5,8.5,11.4,11.3,10.5,11.4,4.3,12.7,24.7,13.3,14.3,11.7,24.4,8.4,9.6,23.3,14,5.6,15.1,16.3,10.2,8,8,9.6,11,16.9,6,32.9,11.87.1"
> x=c(12,19.5,8.5,11.4,11.3,10.5,11.4,4.3,12.7,24.7,13.3,14.3,11.7,24.4,8.4,9.6,23.3,14,5.6,15.1,16.3,10.2,8,8,9.6,11,16.9,6,32.9,11.8,7.1,18,11.8,9.1,7.8,16.5,6.2,11.3,10.6,5.0,15.9,18.1,10.8,13.3,30.1,13.7,8.1,9.7,7.3,13.4)
> mean(x)
[1] 12.81
> sd(x)
[1] 6.063314
> y=c(4,6,4,2,4,6,2,6,3,7,6,5,7,10,3,4,7,7,4,5,5,6,3,2,3,2,5,4,10,9,2,8,4,3,5,5,4,4,2,5,4,7,4,7,6,4,2,5,6,3)
> mean(y)
[1] 4.82
> sd(y)
[1] 2.037305
> z=c(54.52,94.90,26.68,44.73,66.27,67.80,36.04,55.96,70.94,68.73,54.04,48.05,64.16,158.51,84.12,59.20,91.62,126.40,68.45,32.69,78.58,74.43,32.73,48.66,54.66,40.54,34.69,27.91,155.30,120.25,41.20,134.40,37.17,52.09,71.81,59.99,84.17,55.58,39.06,36.48,67.44,60.14,70.38,110.65,104.23,68.17,17.84,103.15,52.15,98.75)
> mean(z)
[1] 68.1282
> sd(z)
[1] 32.34376
>

time pages amtspent
mean 12.81 4.82 68.1282
stan.dev 6.063314 2.037305 32.34376
count 50 50 50
confidence level [10.6013,15.0187} [4.0779,505621] [56.346103,79.910297]